# Well-clipped cones behave themselves under all finite quotients, the cone conjecture under most

### v1

@@ 4.C. Normalizer, centralizer, and symmetries of the invariant cone
Let $\Gamma$ be a group and $G$ be a finite subgroup of $\Gamma$. Then the centralizer $C_{\Gamma}(G)$ is of finite index in $N_{\Gamma}(G)$.
Proof.
Note that $N_{\Gamma}(G)$ acts on $G$ by conjugation, and that $C_{\Gamma}(G)$ is precisely the subgroup of $N_{\Gamma}(G)$ that acts trivially there. This provides an injective group homomorphism
$N_{\Gamma}(G)/C_{\Gamma}(G)\hookrightarrow\mathrm{Aut}(G),$
The order of $\mathrm{Aut}(G)$ being bounded by the finite number $|G|!$ concludes this proof. ∎
Lemma 4.7.
Let $\Gamma_{1}$ and $\Gamma_{2}$ be two commensurable subgroups of a group $\Gamma$, and let $G$ be a finite group of $\Gamma$. Then the centralizers $C_{\Gamma_{1}}(G)$ and $C_{\Gamma_{2}}(G)$ are commensurable.
@@
Consider a group $\Gamma=W\rtimes Q$ and let $G$ be a finite subgroup of $Q$. Then the centralizers satisfy
@@
The next result is the main point of this section. It compares the centralizer subgroup of a finite group $G$ in the symmetry group of a homogeneous cone $\mathscr{A}$ with the symmetries of the invariant cone $\mathscr{A}^{G}$. Recall that the representation $\rho^{G}$ was defined in Definition 4.4.
@@
Let $\mathscr{A}$ be a homogeneous cone in $V=V_{\mathbb{Z}}\otimes\mathbb{R}$. Let $G$ be a finite subgroup of $\mathrm{Aut}(\mathscr{A},V_{\mathbb{Z}})$. Then the inclusion of groups
@@
We denote by $Q_{G}$ the space of corresponding left-cosets. As a quotient of a Lie group by a closed Lie subgroup, $Q_{G}$ is a smooth manifold [22, Theorem 21.10].
We claim that the inclusion
also is of finite index. Indeed, since $G$ is a finite group acting on the rational points of $V$, by Maschke’s theorem, there is a $G$-stable complement $S$, defined over $\mathbb{Q}$, such that $V=V^{G}\oplus S$. Denoting by $S_{\mathbb{Z}}$ the lattice induced by $V_{\mathbb{Z}}$ in $S$, we note that the inclusion $V_{\mathbb{Z}}^{G}\oplus S_{\mathbb{Z}}\subset V_{\mathbb{Z}}$ is of finite index. Computing with block matrices, it is also easy to check that
which, after applying $\rho^{G}$ and using Lemma 4.7, concludes. ∎
We conclude the section with a topological lemma that we just used in the proof of Proposition 4.9.
@@ 4.D. Descent of perfectly clipped cones and more
Before proving a descent result for perfectly clipped cone, we have to address the apparition of an auxiliary self-dual homogeneous cone, a priori smaller than the invariant cone, in the descent result for well-clipped cones given by Proposition 4.5. This is what the discussion of round and simplicial parts prepared us to, in particular the following lemma.
@@
Let $\mathscr{C}$ be a perfectly clipped cone in a self-dual homogeneous cone $\mathscr{A}$ in $V=V_{\mathbb{Z}}\otimes\mathbb{R}$. Let $G<\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$ be a finite subgroup. Then the invariant cone $\mathscr{C}^{G}$ is neatly clipped in a cone $\mathscr{B}$, and we have an inclusion of finite index
@@
Let $\mathscr{C}$ be a perfectly clipped cone in a self-dual homogeneous cone $\mathscr{A}$ in $V=V_{\mathbb{Z}}\otimes\mathbb{R}$. Let $G<\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$ be a finite subgroup. Then the invariant cone $\mathscr{C}^{G}$ is perfectly clipped.
@@ 5. Proof of the main results
The fact that well-clipped cones descend to well-clipped cones is Proposition 4.5. Assume that there exists a rational polyhedral fundamental domain for the action of $\Gamma:=\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$ on the well-clipped cone $\mathscr{C}$. Then by Proposition 3.18, the cone $\mathscr{C}$ is perfectly clipped. By Propositions 4.12 and 3.18, the cone $\mathscr{C}^{G}$ is perfectly clipped in a self-dual homogeneous cone $\mathscr{B}$ such that the inclusion $\rho^{G}(C_{\Gamma}(G)<{\mathrm{Aut}(\mathscr{C}^{G}}^{\circ},\mathscr{B},V_{\mathbb{Z}}^{G})$ is of finite index. Thus, by [25, Proposition 4.6] and Proposition 3.18, the cone ${\mathscr{C}^{G}}^{+}$ admits a rational polyhedral fundamental domain for the action of $\rho^{G}(C_{\Gamma}(G))$. ∎
@@
Let $G$ be a finite subgroup of $\mathrm{Aut}(X,\Delta)$. Denoting by $p:X\to Y$ the quotient map, note that $p^{*}$ is injective and
$N^{1}(X)^{G}=p^{*}N^{1}(X/G),\quad\mathrm{Mov}(X)\cap N^{1}(X)^{G}=\quad p^{*}\mathrm{Mov}(X/G),$
$\rho^{G}(N_{\mathrm{PsAut}^{*}(X,\Delta)}(G))<\quad\mathrm{PsAut}^{*}(X/G,\Delta_{G}),$
where $\mathrm{PsAut}^{*}$ denotes the image of the pseudoautomorphism group $\mathrm{PsAut}$ in $\mathrm{GL}(N^{1}(\cdot))$, and $(X/G,\Delta_{G})$ is the quotient pair of $(X,\Delta)$ by $G$. Recall that $\mathrm{PsAut}^{*}(X,\Delta)$ is contained in $\Gamma:=\mathrm{Aut}(\mathrm{Mov}(X)^{\circ},\mathscr{A},N^{1}(X)_{\mathbb{Z}-{\rm Weil}})$, where $\mathscr{A}$ denotes a self-dual homogeneous cone in which $\mathrm{Mov}(X)$ is well-clipped.
Since $(X,\Delta)$ satisfies the movable cone conjecture, there is a rational polyhedral fundamental domain for the action of $\Gamma$ on $\mathrm{Mov}^{+}(X)$. Thus, by [25, Proposition 4.6], the groups $\mathrm{PsAut}^{*}(X,\Delta)$ and $\Gamma$ are commensurable. By Theorem 1.7, there is a rational polyhedral fundamental domain for the action of $\rho^{G}(C_{\Gamma}(G))$ on $\mathrm{Mov}^{+}(X/G)$. By Lemmas 4.6, 4.7 and [25, Proposition 4.6], the cone $\mathrm{Mov}^{+}(X/G)$ thus admits a rational polyhedral fundamental domain for the action of $\rho^{G}(N_{\mathrm{PsAut}^{*}(X,\Delta)}(G))$, and a fortiori [25, Proposition 4.1, Application 4.14] for the larger group $\mathrm{PsAut}^{*}(X/G,\Delta_{G})$. ∎
@@
Recall that $X=A\times Y\times\prod_{j=1}^{s}S_{j},$ where $A$ is an abelian variety, $Y$ is a product of primitive symplectic varieties with canonical singularities, and each $S_{j}$ is a smooth rational surface underlying a klt Calabi–Yau pair $(S_{j},\Delta_{j})$. Let $G$ be a finite subgroup of $\mathrm{Aut}(X,\Delta)$. By [9, Lemma 4.6] the $G$-action preserves the decomposition, up to permutation of potential isomorphic factors.
@@
By [39, Corollary 4.4.4], each $Y_{k}$ admits a $G\cap\mathrm{Aut}(Y_{k})^{n_{k}}$-equivariant $\mathbb{Q}$-factorial terminalization $\hat{Y_{k}}$. In particular,
@@
We now prove the movable cone conjecture for any quotient pair of $(X,\Delta)$. By [15, Exercise III.12.6] and Lemma 2.16, the movable cone of $X$ is the direct sum of the movable cones of its factors. Thus and by [45, 38, 23], the pair $(X,\Delta)$ satisfies the movable cone conjecture. We checked in Example 3.6, see in particular Items (1), (4), (6), (7) that the movable cone of each factor of $X$ is well clipped, and in fact
@@
where $W$ is the set of pullbacks of primitive wall divisors of the primitive symplectic terminal $\mathbb{Q}$-factorial factors (see [23, Definition 7.1]). By [23, Proposition 7.7], the set of squares $\{q(w)\mid w\in W\}$ is bounded and contained in $\mathbb{Z}_{<0}$. Restricting to the $G$-invariant Néron-Severi subspace, we still have
@@
The elements of $W_{G}$ clearly remain of bounded negative squares, thus by [30, Proposition 3.4], for any rational polyhedral cone $\Pi$, only finitely many of their orthogonal hyperplanes intersect $\Pi^{\circ}$. This checks the assumption of Lemma 2.19, and applying it to the pair $(X/G,\Delta_{G})$ concludes the proof. ∎

### v2

@@ 4.C. Centralizer and symmetries of the invariant cone
Let $\Gamma_{1}$ and $\Gamma_{2}$ be two commensurable subgroups of a group $\Gamma$, and let $G$ be a subgroup of $\Gamma$. Then the centralizers $C_{\Gamma_{1}}(G)$ and $C_{\Gamma_{2}}(G)$ are commensurable.
@@
Consider a group $\Gamma=W\rtimes Q$ and let $G$ be a subgroup of $Q$. Then the centralizers satisfy
@@
The next result is the main point of this section. It compares the centralizer subgroup of a group $G$ acting with finite orbits on a homogeneous cone $\mathscr{A}$ with the symmetries of the invariant cone $\mathscr{A}^{G}$. Recall that the representation $\rho^{G}$ was defined in Definition 4.4.
@@
Let $\mathscr{A}$ be a homogeneous cone in $V=V_{\mathbb{Z}}\otimes\mathbb{R}$. Let $G$ be a subgroup of $\mathrm{Aut}(\mathscr{A},V_{\mathbb{Z}})$ acting with finite orbits. Then the inclusion of groups
@@
Before proving this result, we establish two preparatory lemmas. The first one is a spin-off on Maschke’s theorem.
Lemma 4.9.
Let $V$ be a finite dimensional $\mathbb{R}$-vector space and $G<{\rm GL}(V)$ be a subgroup acting with finite orbits. Then $V^{G}$ admits a $G$-invariant complement $S\subset V$.
@@
Let $p:V\twoheadrightarrow V^{G}$ be any linear projection. We define
$\overline{p}:v\in V\mapsto\frac{1}{|G\cdot v|}\sum_{w\in G\cdot v}p(w)\in V^{G}.$
Clearly, $\overline{p}$ restricts to the identity on $V^{G}$. Moreover, for all $g\in G$, we have $\overline{p}(g(v))=\overline{p}(v)$. Finally, we check that $\overline{p}$ is linear: Since $G$ acts with finite orbits, we have a subgroup inclusion of finite index:
$H:={\rm Stab}(v_{1})\cap{\rm Stab}(v_{2})\quad<\quad{\rm Stab}(v_{1}+v_{2})$
for any $v_{1},v_{2}\in V$. This implies:
$\overline{p}(v_{1}+v_{2})=\frac{|H|}{|G|}\sum_{g\in G/H}p(g(v_{1}+v_{2}))=\overline{p}(v_{1})+\overline{p}(v_{2}),$
as wished. Therefore, $\overline{p}$ is a $G$-equivariant projection from $V$ to $V^{G}$, and its kernel is a $G$-invariant complement to $V^{G}$. ∎
The following lemma is topological.
@@
We can now prove Proposition 4.8.
Proof of Proposition 4.8.
We denote by $Q_{G}$ the space of corresponding left-cosets. As a quotient of a Lie group by a closed Lie subgroup, $Q_{G}$ is a smooth manifold [24, Theorem 21.10].
We claim that the inclusion
also is of finite index. Indeed, since $G$ is a group acting with finite orbits on the rational points of $V$, by Maschke’s theorem (see Lemma 4.9), there is a $G$-stable complement $S$, defined over $\mathbb{Q}$, such that $V=V^{G}\oplus S$. Denoting by $S_{\mathbb{Z}}$ the lattice induced by $V_{\mathbb{Z}}$ in $S$, we note that the inclusion $V_{\mathbb{Z}}^{G}\oplus S_{\mathbb{Z}}\subset V_{\mathbb{Z}}$ is of finite index. Computing with block matrices, it is also easy to check that
which, after applying $\rho^{G}$ and using Lemma 4.6, concludes. ∎
@@ 4.D. Descent of perfectly clipped cones and more
Before proving a descent result for perfectly clipped cones, we have to address the apparition of an auxiliary self-dual homogeneous cone, a priori smaller than the invariant cone, in the descent result for well-clipped cones given by Proposition 4.5. This is what the discussion of round and simplicial parts prepared us to, in particular the following lemma.
@@
Let $\mathscr{C}$ be a perfectly clipped cone in a self-dual homogeneous cone $\mathscr{A}$ in $V=V_{\mathbb{Z}}\otimes\mathbb{R}$. Let $G<\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$ be a subgroup acting with finite orbits. Then the invariant cone $\mathscr{C}^{G}$ is neatly clipped in a cone $\mathscr{B}$, and we have an inclusion of finite index
@@
Let $\mathscr{C}$ be a perfectly clipped cone in a self-dual homogeneous cone $\mathscr{A}$ in $V=V_{\mathbb{Z}}\otimes\mathbb{R}$. Let $G<\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$ be a subgroup acting with finite orbits. Then the invariant cone $\mathscr{C}^{G}$ is perfectly clipped.
@@ 5. Proof of the main results
The fact that well-clipped cones descend to well-clipped cones is Proposition 4.5. Assume that there exists a rational polyhedral fundamental domain for the action of $\Gamma:=\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$ on the well-clipped cone $\mathscr{C}$. Then by Proposition 3.18, the cone $\mathscr{C}$ is perfectly clipped. By Propositions 4.12 and 3.18, the cone $\mathscr{C}^{G}$ is perfectly clipped in a self-dual homogeneous cone $\mathscr{B}$ such that the inclusion $\rho^{G}(C_{\Gamma}(G))<{\mathrm{Aut}(\mathscr{C}^{G}}^{\circ},\mathscr{B},V_{\mathbb{Z}}^{G})$ is of finite index. Thus, by [27, Proposition 4.6] and Proposition 3.18, the cone $\mathscr{C}^{G}$$^{+}$ admits a rational polyhedral fundamental domain for the action of $\rho^{G}(C_{\Gamma}(G))$. ∎
@@
Let $G$ be a finite subgroup of $\mathrm{Aut}(X,\Delta)$ and let $\overline{G}$ be the group $G\times{\rm Gal}(\overline{k}/k)$ acting on the base change to the algebraic closure $(\overline{X},\overline{\Delta})$. Note that its induced action on the Néron–Severi space $N^{1}(\overline{X})_{\mathbb{R}}$ has finite orbits by [23, Paragraphs 16.1, 16.2]. Denoting by $p:\overline{X}\to Y$ the quotient map by the action of $\overline{G}$, note that $p^{*}$ is injective and
$N^{1}(\overline{X})^{G}=p^{*}N^{1}(Y),\quad\mathrm{Mov}(\overline{X})\cap N^{1}(\overline{X})^{G}=\quad p^{*}\mathrm{Mov}(Y),$
$\rho^{G}(C_{\mathrm{PsAut}^{*}(\overline{X},\overline{\Delta})}(G))<\quad\mathrm{PsAut}^{*}(Y,\Delta_{G}),$
where $\mathrm{PsAut}^{*}$ denotes the image of the pseudoautomorphism group $\mathrm{PsAut}$ in $\mathrm{GL}(N^{1}(\cdot))$, and $(Y,\Delta_{G})$ is the quotient pair of $(X,\Delta)$ by $G$. Recall that $\mathrm{PsAut}^{*}(\overline{X},\overline{\Delta})$ is contained in $\Gamma:=\mathrm{Aut}(\mathrm{Mov}(\overline{X})^{\circ},\mathscr{A},N^{1}(\overline{X})_{\mathbb{Z}-{\rm Weil}})$, where $\mathscr{A}$ denotes a self-dual homogeneous cone in which $\mathrm{Mov}(\overline{X})$ is well-clipped.
Since $(\overline{X},\overline{\Delta})$ satisfies the movable cone conjecture, there is a rational polyhedral fundamental domain for the action of $\Gamma$ on $\mathrm{Mov}^{+}(\overline{X})$. Thus, by [27, Proposition 4.6], the groups $\mathrm{PsAut}^{*}(\overline{X},\overline{\Delta})$ and $\Gamma$ are commensurable. By Theorem 1.8, there is a rational polyhedral fundamental domain for the action of $\rho^{G}(C_{\Gamma}(G))$ on $\mathrm{Mov}^{+}(Y)$. By Lemmas 4.6 and [27, Proposition 4.6], the cone $\mathrm{Mov}^{+}(Y)$ thus admits a rational polyhedral fundamental domain for the action of $\rho^{G}(C_{\mathrm{PsAut}^{*}(\overline{X},\overline{\Delta})}(G))$, and a fortiori [27, Proposition 4.1, Application 4.14] for the larger group $\mathrm{PsAut}^{*}(Y,\Delta_{G})$. ∎
Remark 5.1.
In principle, one could use that $\mathrm{PsAut}^{*}(X/G,\Delta_{G})$ entails the larger normalizer subgroup $\rho^{G}(N_{\mathrm{PsAut}^{*}(\overline{X},\overline{\Delta})}(G))$. However, working with the centralizer subgroup is somewhat simpler.
Proof of Corollary 1.7.
Let $X$ be a torsor over an abelian variety over a perfect field $k$. The base change $\overline{X}$ to the algebraic closure of $k$ is an abelian variety, thus by the work of Prendergast–Smith in [41] has a self-dual homogeneous movable cone and satisfies the movable cone conjecture. Theorem 1.6 concludes. ∎
@@
Recall that $X=A\times Y\times\prod_{j=1}^{s}S_{j},$ where $A$ is an abelian variety, $Y$ is a product of primitive symplectic varieties with canonical singularities, and each $S_{j}$ is a smooth rational surface underlying a klt Calabi–Yau pair $(S_{j},\Delta_{j})$. Let $G$ be a finite subgroup of $\mathrm{Aut}(X,\Delta)$. By [10, Lemma 4.6] the $G$-action preserves the decomposition, up to permutation of potential isomorphic factors.
@@
By [42, Corollary 4.4.4], each $Y_{k}$ admits a $G\cap\mathrm{Aut}(Y_{k})^{n_{k}}$-equivariant $\mathbb{Q}$-factorial terminalization $\hat{Y_{k}}$. In particular,
@@
We now prove the movable cone conjecture for any quotient pair of $(X,\Delta)$. By [16, Exercise III.12.6] and Lemma 2.16, the movable cone of $X$ is the direct sum of the movable cones of its factors. Thus and by [48, 41, 25], the pair $(X,\Delta)$ satisfies the movable cone conjecture. We checked in Example 3.6, see in particular Items (1), (4), (6), (7) that the movable cone of each factor of $X$ is well clipped, and in fact
@@
where $W$ is the set of pullbacks of primitive wall divisors of the primitive symplectic terminal $\mathbb{Q}$-factorial factors (see [25, Definition 7.1]). By [25, Proposition 7.7], the set of squares $\{q(w)\mid w\in W\}$ is bounded and contained in $\mathbb{Z}_{<0}$. Restricting to the $G$-invariant Néron-Severi subspace, we still have
@@
The elements of $W_{G}$ clearly remain of bounded negative squares, thus by [32, Proposition 3.4], for any rational polyhedral cone $\Pi$, only finitely many of their orthogonal hyperplanes intersect $\Pi^{\circ}$. This checks the assumption of Lemma 2.19, and applying it to the pair $(X/G,\Delta_{G})$ concludes the proof. ∎